the uptown parking garage charges $6.25 for the first hour and $0.60 for each additional hour. the downtown parking garage charges $4.75 for the first hour and $0.90 for each additional hour. for how many hours will the two garages charge an equal amount of money?
Answer by macston(5194)

.
H=additional hours; H+1=number of hours when cost is the same
.
$6.25+$0.60H=$4.75+$0.90H
$1.50=$0.30H
5=H
.
H+1=5+1=6
.
ANSWER: The cost is the same after a total of 6 hours.